### v1

# STCast: Adaptive Boundary Alignment for Global and Regional Weather Forecasting

Abstract
To gain finer regional forecasts, many works have explored the regional integration from the global atmosphere, e.g., by solving boundary equations in physics-based methods or cropping regions from global forecasts in data-driven methods. However, the effectiveness of these methods is often constrained by static and imprecise regional boundaries, resulting in poor generalization ability. To address this issue, we propose Spatial-Temporal Weather Forecasting (STCast), a novel AI-driven framework for adaptive regional boundary optimization and dynamic monthly forecast allocation. Specifically, our approach employs a Spatial-Aligned Attention (SAA) mechanism, which aligns global and regional spatial distributions to initialize boundaries and adaptively refines them based on attention-derived alignment patterns. Furthermore, we design a Temporal Mixture-of-Experts (TMoE) module, where atmospheric variables from distinct months are dynamically routed to specialized experts using a discrete Gaussian distribution, enhancing the model’s ability to capture temporal patterns. Beyond global and regional forecasting, we evaluate our STCast on extreme event prediction and ensemble forecasting. Experimental results demonstrate consistent superiority over state-of-the-art methods across all four tasks.
Introduction
Why need global forecasts to support regional forecasting? Achieving accurate, kilometre-scale regional weather forecasting is still a formidable scientific task with far-reaching socio-economic impact. Existing strategies typically fall into two paths: training a dedicated regional model or extracting the regional prediction from a global forecast. Traditional Numerical Weather Prediction (NWP) (Bauer, Thorpe, and Brunet 2015; Lynch 2008; Kalnay 2002) methods solve partial differential equations (PDEs) at finer resolutions, but incur prohibitively high computational cost. Recent data-driven approaches (Bi et al. 2023; Chen et al. 2023b; Lam et al. 2023; Bodnar et al. 2025; Subich et al. 2025; Wu et al. 2025) significantly reduce this cost by neural networks. However, these models often rely on patch embeddings that downsample input variables, resulting in the loss of fine-grained local details. That is, training a global model at 1 km resolution (approximately 0.01°, or ) would be computationally infeasible. Conversely, restricting training to a high-resolution region neglects cross-regional dependencies that are critical for accurate forecasting. Thus, both direct training of high-resolution regional models and extracting them from high-resolution global forecasts are impractical. These limitations highlight the need for hybrid frameworks that couple low-resolution global forecasts with high-resolution regional forecasting in a computationally efficient manner.
Why adaptive boundaries must cover the global area? While hybrid global–regional frameworks are gaining attention, existing coupling strategies, whether based on traditional NWP (Lundquist, Chow, and Lundquist 2010; Mani 2012) or AI models (Gao et al. 2025; Adamov et al. 2025; Nipen et al. 2024), typically define regional boundaries using only adjacent areas. This local perspective contradicts the well-established Atmosphere–Ocean–Land–Biosphere Coupling Theory (Manabe and Bryan 1969; Zhang, Tian, and Wang 2018), which posits that any points in the regional atmosphere are influenced by the entire Earth system. For example, Siberian cold surges can trigger East-Asian cold waves, and surface heating over the Tibetan Plateau can simultaneously alter East Asian monsoons and North American jet stream (Wu et al. 2023). Thus, the true boundary for a region is not its neighbors, but the entire Earth.
To address these challenges, we introduce STCast, a Spatial–Temporal Forecasting framework that explicitly models the evolving global–regional correlations within Earth system. Unlike prior methods that restrict the boundaries to neighboring regions, STCast initializes the global-regional distributions using spatial-aligned attention (SAA) and continuously refines them during training. Beyond spatial boundaries modeling, STCast further captures temporal variability by routing monthly atmospheric inputs to specialized experts via Temporal Mixture-of-Experts (TMoE), using a discrete Gaussian distribution. Together, SAA and TMoE enable STCast to deliver accurate and generalizable regional forecasts by incorporating both global spatial influences and fine-grained temporal patterns.
Spatial-Aligned Attention (SAA) incorporates a learnable global-regional distribution into linear cross-attention, enabling adaptive aggregation of global atmospheric information for regional forecasting. To couple the global and regional variables, SAA employs two key mechanisms: (1) a Manhattan distance metric to quantify spatial separation from the target region, and (2) an exponential distance-decay function to initialize the learnable global-regional distribution, ensuring weaker influence from distant regions. This prior modulates the attention weights by element-wise multiplication and is further refined during training. As a result, the global–regional correlation evolves dynamically, aligning spatial dependencies with physical intuition throughout the optimization process.
Temporal Mixture-of-Experts (TMoE) enhances the standard MoE framework by integrating a month-specific Gaussian prior to guide expert routing. It operates through three key mechanisms: (1) Learning a Gaussian distribution for each month to represent its temporal characteristics; (2) Modulating expert routing weights with this distribution, ensuring that weights decay with increasing temporal distance from an expert; (3) Enabling multi-expert activation to enhance routing diversity. This design facilitates dynamic input-to-expert assignment while preserving temporal specialization and improving generalization across time.
Regional Forecasting Experiments. As illustrated in Fig. 1.(1), we compare three regional weather forecasting strategies, including previous AI methods (Fig. 1a), directly training on the target region (Fig. 1b), and our proposed STCast (Fig. 1c). Unlike existing approaches that statically concatenate adjacent areas to the target region, STCast establishes a learnable global-regional distribution to adaptively aggregate low-resolution global forecasts into high-resolution regions. Quantitative results in Fig. 1.(2) demonstrate that STCast achieves the best performance across all variables in terms of both Mean RMSE and ACC, outperforming Direct Train and OneForecast. These results validate the effectiveness of our dynamic, Earth-aware boundary mechanism over static neighbor-based coupling.
In conclusion, the contributions of this work include:
We propose an AI-based method to extract adaptive regional boundary from our Spatial-Aligned Attention (SAA) module. The approach is initialized with global-regional distribution and optimized during training.
We introduce the Spatial-Temporal Forecasting Framework (STCast) for weather forecasting, featuring a novel Temporal Mixture-of-Experts (TMoE) architecture. This component dynamically allocates forecasting tasks across different months to specialized expert models, enhancing temporal adaptability.
Extensive experiments across four critical weather forecasting tasks, including low-resolution global forecasts, high-resolution regional forecasts, typhoon track prediction, and ensemble forecasting, demonstrate that STCast achieves state-of-the-art performance, significantly outperforming existing methods.
Related work
Global-Regional Weather Coupling
Accurate global–regional coupling remains a core challenge in regional forecasting due to the difficulty of boundary specification. NWP models address this by solving PDEs under prescribed boundary, e.g., sponge layers (Mani 2012), Dirichlet (Hidayatullah et al. 2019), and Neumann conditions (Sabathier et al. 2023). In contrast, AI approaches (Gao et al. 2025; Adamov et al. 2025; Nipen et al. 2024) bypass boundary equations by concatenating a fixed neighborhood in global area to region, resulting in static and local coupling.
The recent work, OneForecast (Gao et al. 2025), tackles the same 4 tasks as ours by concatenating neighboring low-resolution global forecasts with high-resolution regional variables. In contrast, our method replaces static concatenation with a transformer-based framework that adaptively models global–regional correlations, guided by a learned prior. This enables dynamic boundary refinement and long-range dependency modeling beyond local neighborhoods.
Data-Driven Weather Forecasting
Prior to deep learning, numerical weather prediction (NWP) prevailed, producing forecasts by solving PDEs on high-resolution global grids (Bauer, Thorpe, and Brunet 2015; Lynch 2008; Kalnay 2002). These methods deliver physically analysis and rigorously validated forecasts (Molteni et al. 1996; Ritchie et al. 1995), but their computation at inference remains prohibitively high, especially at fine scales.
The emergence of large-scale atmospheric reanalyses has catalysed a shift toward data-driven forecasting. Early exemplars, FourCastNet (Kurth et al. 2023) and Pangu-Weather (Bi et al. 2023), employ Fourier Neural Operators (FNO) (Li et al. 2021) and 3D Swin Transformers (Liu et al. 2021), respectively, to approximate atmospheric evolution. Subsequent research has bifurcated into two streams: (i) neural operators, such as KNO (Xiong et al. 2023) and SFNO (Bonev et al. 2023), that directly learn the temporal evolution operator; and (ii) neural networks, including FengWu (Chen et al. 2023a), FengWu-ghr (Han et al. 2024), GraphCast (Lam et al. 2023), FuXi (Chen et al. 2023b), GenCast (Price et al. 2025), and Stormer (Nguyen et al. 2025), which leverage inductive biases tailored to atmosphere. All achieve competitive skill with limited computation.
Inspired by the success of MoE in LLMs (Shazeer et al. 2017), recent studies have begun to integrate MoE into weather forecasting. VAMoE (Chen et al. 2025) extends MoE to incremental weather forecasting; EWMoE (Gan et al. 2025) augments FourCastNet with MoE layers. In contrast, we propose TMoE that explicitly partitions atmospheric inputs by month and dynamically routes them to specialized temporal experts. This structure enables STMoE to capture inter-month variability and intra-month correlation.
Methodology
We give a unified framework for 4 tasks: low-resolution global prediction, high-resolution regional prediction, typhoon track forecasting, and ensemble forecasting. Section Overview formally defines each task. We then introduce Spatial-Aligned Attention (SAA) that fuses global and regional atmospheric variables via a learnable global-regional distribution, and Temporal Mixture-of-Experts (TMoE) that allocates monthly data to specialized experts.
Overview
For the weather forecasting task, the AI model infers future atmospheric states from historical fields , i.e., . Here, comprises upper-air variables on 13 pressure levels and surface variables , where and denote the number of variables per pressure and surface level, respectively.
As illustrated in Fig. 2, a framework unifies four subtasks: global deterministic forecasting , high-resolution regional forecasting , typhoon track prediction , and ensemble forecasting . In Fig. 2(a), we introduce Temporal Mixture-of-Experts (TMoE) and further integrate Flash-Attention (Dao et al. 2022) with MoE (Shazeer et al. 2017). Recognizing the pronounced seasonal variability of atmospheric states, TMoE treats monthly forecast as a distinct tasks and assigns it to multiple dedicated experts. Global forecasting is expressed as , where denotes the global variables. For regional forecasting, we employ a global–regional coupling strategy in Fig. 2(b). A Spatial-Aligned Attention (SAA) module fuses global variables with regional inputs to yield high-resolution predictions: . The predicted MSL is subsequently used to infer typhoon tracks in Fig. 2(c). Beyond deterministic forecasts, we evaluate probabilistic skill for both long-range and ensemble scenarios in Fig. 2(d). Perlin noise is injected into initial state , and the model is run times; the ensemble mean provides probabilistic forecast.
The principal contributions of this work are the TMoE and SAA modules, detailed in the following subsections.
Spatial-Aligned Attention
As shown in Fig. 3, the Spatial-Aligned Attention (SAA) module employs global features as Query and Key, while utilizing regional features as Value. Unlike previous approaches that rely on static boundaries, our SAA module dynamically couples global and regional features through linear cross-attention at each block. This innovative design learns global-regional distribution from attention maps while maintaining computational efficiency through linear attention mechanisms, which effectively reduce the processing overhead.
For precise quantification of spatial relationships, SAA calculates Manhattan distance between each global point and target region. This distance metric is defined as:
where denotes the center coordinates of target region, and represent its height and width. This efficient formulation introduces negligible overhead.
Next, the global–regional prior is derived from an exponential distance-decay function that monotonically reduces correlation as distance increases. The function is:
where exp and denote the base of the natural logarithm and the decay factor, respectively.
SAA establishes an optimal distribution by computing the Hadamard product between the initial global-regional distribution and the attention map. This trainable prior distribution serves a dual purpose: it guides the optimization process while being progressively refined, capturing the spatial relationships and learning latent correlations between global and regional atmospheric patterns.
Model Metric 6-hour 1-day 4-day 7-day 10-day RMSE ACC RMSE ACC RMSE ACC RMSE ACC RMSE ACC Pangu-weather(Bi et al. 2023) 0.0826 0.9876 0.1571 0.9581 0.3380 0.8167 0.5092 0.5738 0.6215 0.3542 Graphcast(Lam et al. 2023) 0.0626 0.9928 0.1304 0.9705 0.2861 0.8705 0.4597 0.6692 0.6009 0.4275 Fuxi(Chen et al. 2023b) 0.0987 0.9820 0.1708 0.9511 0.4128 0.7379 0.5972 0.4446 0.6981 0.2391 Oneforecast(Gao et al. 2025) 0.0549 0.9943 0.1231 0.9737 0.2732 0.8825 0.4468 0.6888 0.5918 0.4457 Ours 0.0617 0.9956 0.1197 0.9740 0.2578 0.8927 0.4348 0.7019 0.5763 0.4715
Temporal Mixture-of-Experts
Acknowledging the discrepancy of atmospheric variables across different months, the Temporal Mixture-of-Experts (TMoE) framework treats forecasting for each month as relatively independent tasks and organizes these tasks using the Mixture-of-Experts (MoE). To assign training tasks for different months to specialized experts, TMoE employs a rotating discrete Gaussian distribution that directs the experts in training atmospheric variables across various months. The peak of the Gaussian distribution is rotated to correspond with the month of the input variables. The discrete Gaussian distribution is defined as follows:
where and are the mean and variance of the distribution, and is a discrete series, , denoting 12 months in one year. To fit the atmospheric dataset, those two hyper-parameters are set to learnable during training.
Following the discrete Gaussian series, we perform rotational alignment of the month series to correspond with input variables. This alignment ensures a monotonic decrease in activation probability as temporal distance from the target month increases. Through this mechanism, input variables become distinguishable by their month. The aligned month series is subsequently encoded into continuous embedding representations via a MLP. These temporal embeddings serve as latent features that inform and optimize the expert selection process within TMoE.
In TMoE, the gating network first derives a weight vector and an index tensor from the input variables . Month-specific information is incorporated by concatenating this index with the 12-dimensional month embedding . The resulting feature is then fed into a softmax layer that selects the Top-K experts. These experts are subsequently activated to model the conditional distribution of inputs for current month. The entire procedure is formulated as:
where I denotes index, which selects Top-K experts.
Compared to prior MoE methods that employ implicit expert allocation strategies with auxiliary losses, TMoE introduces an explicit month embedding mechanism to assign input variables to specialized experts with limited computation. This explicit guidance more effectively prevents MoE homogenization during training.
Ablation Studies Metric 6-hour 1-day 4-day 7-day 10-day RMSE ACC RMSE ACC RMSE ACC RMSE ACC RMSE ACC High-resolution Regional Forecasts w/o SAA 0.0767 0.9762 0.1802 0.8675 0.4001 0.4127 0.5297 0.4229 0.7610 0.2541 w/o Global-Regional Distribution 0.0694 0.9805 0.1631 0.8864 0.3794 0.6718 0.5082 0.4566 0.7192 0.3286 w SAA 0.0493 0.9946 0.0854 0.9854 0.2068 0.9203 0.3712 0.7442 0.4945 0.5433 Low-resolution Global Forecasts w/o TMoE 0.0751 0.9915 0.1451 0.9714 0.3249 0.8201 0.5109 0.5412 0.6426 0.3184 w/o Month Embedding 0.0744 0.9928 0.1346 0.9764 0.2865 0.8180 0.4631 0.5941 0.6049 0.3559 w TMoE 0.0617 0.9956 0.1197 0.9740 0.2578 0.8927 0.4348 0.7019 0.5763 0.4715
2024.11.04, T+0
2024.11.06, T+12
2024.11.06, T+12
| Model | Forecast Day | |||
| 7-day | 8-day | 9-day | 10-day | |
| Pangu | 0.4875 | 0.5321 | 0.5742 | 0.6213 |
| Pangu (ENS) | 0.4435 | 0.4743 | 0.4979 | 0.5205 |
| Graphcast | 0.4440 | 0.4923 | 0.5346 | 0.5823 |
| Graphcast (ENS) | 0.4412 | 0.4759 | 0.5072 | 0.5331 |
| Fuxi | 0.5928 | 0.6314 | 0.6604 | 0.6968 |
| Fuxi (ENS) | 0.4898 | 0.5175 | 0.5353 | 0.5498 |
| OneForecast | 0.4268 | 0.4834 | 0.5313 | 0.5809 |
| OneForecast (ENS) | 0.4393 | 0.4699 | 0.4951 | 0.5167 |
| Ours | 0.3892 | 0.4285 | 0.4708 | 0.5107 |
| Ours (ENS) | 0.3893 | 0.4284 | 0.4713 | 0.5113 |
Experiments
Main Results
Low-resolution Global Weather Forecasting. We evaluate forecasting performance using two standard metrics: RMSE and ACC. Due to significant scale variations across atmospheric variables, direct comparison using absolute values is infeasible. We therefore present normalized RMSE and ACC scores in Tab. 1, where STCast demonstrates consistent superiority over baselines across all benchmarks, with particularly significant gains in long-term predictions. Further validation through real-value RMSE and ACC comparisons (1-10 day) in Fig. 5 confirms STCast’s state-of-the-art performance across multiple variables. This enhancement is attributed to our month-specific training strategy, which effectively captures both seasonal dependencies and month-to-month variations in weather systems. Complementary visualization in Fig. 6 compares spatial error distributions for five key variables across three methods, providing qualitative evidence of STCast’s reduced prediction uncertainty.
High-resolution Regional Weather Forecasting. As demonstrated in Fig. 1, we compare mean RMSE and ACC scores across five surface variables against Direct Training and OneForecast (Gao et al. 2025). Quantitative analysis reveals that direct-trained STCast (without dynamic boundary) and OneForecast achieve comparable performance. However, implementing our dynamic boundary condition in STCast yields significant improvements: mean RMSE decreases by 0.05 while mean ACC increases by 0.1. This enhancement confirms the critical role of adaptive boundary modeling in regional forecasting systems. Complementary visualization of 6-hour regional forecasts for U10 and MSL in Fig. 8 provides performance validation. Error analysis demonstrates STCast’s superior accuracy, achieving near-zero relative errors of just 0.7% for U10 and 0.1% for MSL - substantially lower than competitors.
Long-term and Ensemble Weather Forecasting. As illustrated in Fig. 7, we compare 100-day Z500 predictions from top-performing models. GraphCast exhibits significant predictive degradation in high-latitude regions, while OneForecast maintains functionality but shows substantial poleward deviations from ground truth. In contrast, STCast demonstrates consistent alignment with observations across all latitudes, despite minor localized discrepancies. These long-term forecasts confirm STCast’s superior performance in high-latitude prediction tasks. We further evaluate ensemble forecasting capabilities using 10 initial conditions (commencing 00:00 UTC 1 January 2020 at 12-hour intervals) in Tab. 3. Quantitative assessment via normalized Mean RMSE demonstrates that both STCast and its ensemble variant significantly outperform four competing methods, with our ensemble approach achieving the lowest error distribution across all initialization times.
Extreme Events Assessment. Extreme weather events, particularly tropical cyclones, pose significant societal risks, necessitating accurate prediction capabilities (Wang et al. 2025a, b). To evaluate our method’s performance under such critical conditions, we analyze two recent typhoon events: Typhoon Ewiniar (May 2024) and Typhoon Yinxing (November 2024) (Ying et al. 2014; Lu et al. 2021). As visualized in Fig. 9d-e, STCast’s 72-hour track forecasts demonstrate substantially closer alignment with observed paths compared to operational ECMWF predictions for both systems. This improved track accuracy highlights STCast’s enhanced capability for extreme event forecasting. Complementary visualization in Fig. 9a-c further dissects Typhoon Yinxing’s evolution, contrasting initial conditions, STCast predictions, and ground-truth. Our model consistently captures the cyclone’s structural development and translational dynamics, validating its physical representation of intense meteorological systems. More numerical comparison is shown in Appendix.
More Competitors. ClimaX (Nguyen et al. 2023), EWMoE (Gan et al. 2025), Keisler (Keisler 2022), Stormer (Nguyen et al. 2025), VAMoE (Chen et al. 2025), FourCastNet (Kurth et al. 2023), ClimODE (Verma, Heinonen, and Garg 2024), WeatherGFT (Xu et al. 2024), and GenCast (Price et al. 2025) are shown in Appendix.
Computation comparisons, additional visualization, and additional results are provided in Appendix.
Ablation Study
To further verify the effectiveness of each proposed module and strategy, we conduct comprehensive ablation studies reported in Tab. 2. The experiments are split into two groups. High-resolution regional forecasting: (1) STCast w/o SAA: we remove the SAA module and follow the same protocol as Oneforecast to predict regional variables; (2) STCast w/o Global-Regional Distribution: we discard the global-regional distribution initialization in SAA and instead use random initialization for the global-regional correlation; (3) STCast w SAA: the complete STCast model. Low-resolution global forecasting: (4) STCast w/o TMoE: we replace the Temporal Mixture-of-Experts (TMoE) with MLP block; (5) STCast w/o Month Embedding: we remove the month embedding from TMoE and fall back to a classical Mixture-of-Experts; (6) STCast w TMoE: the complete STCast model. Comparisons among (1)–(3) and (4)–(6) reveal that removing any component consistently degrades performance on both regional and global tasks. While the absence of SAA or TMoE causes noticeable drops, the most substantial drops occur when eliminating global-regional distribution (regional: +0.22 RMSE in 10-day) and month embedding (global: +0.13 RMSE in 10-day). These results confirm the critical roles of every component and setting in enhancing the overall effectiveness of STCast.
Conclusion
In this work, we introduce an adaptive attention map within the Spatial-Aligned Attention (SAA) module to provide dynamic boundary conditions for regional forecasting. Beyond regional task, we embed a Temporal Mixture-of-Experts (TMoE) into the Spatial-Temporal Forecasting (STCast), casting weather prediction as a multi-task problem and delegating monthly sub-tasks to specialized experts. Consequently, STCast simultaneously addresses 4 distinct challenges: low-resolution global forecasting, high-resolution regional forecasting, extreme-event assessment, and ensemble weather forecasting. Comprehensive experiments and ablation studies confirm that STCast consistently outperforms competing methods across all evaluated scenarios.
Appendix
STCast: Adaptive Boundary Alignment for Global and Regional
Weather Forecasting
Dataset Details
Dataset and Baselines
In this work, we conduct experiments on a popular weather dataset, i.e., ERA5111https://cds.climate.copernicus.eu/ (Hersbach et al. 2020), provided by the ECMWF (Molteni et al. 1996). As shown in Tab. 4, ERA5 dataset is a reanalysis atmospheric dataset, consisting of the atmospheric variables from 1979 to the present day with a 0.25°spatial resolution with . The atmospheric variables include 5 upper-air variables (Z, Q, U, V, T) on 13 levels and 5 surface variables (T2M, U10, U10, MSL, SP). The low-resolution global forecasting is trained on 70 variables () with 40 years atmospheric dataset from 1979 to 2019 with 1.4°spatial resolution. At the same time, the high-resolution regional weather forecasting task is trained in the same period with 5 surface variables on Eastern Asia (7.5°W114°E-36°W172.5°E) with 0.25°spatial resolution.
Data processing
To address disparities among variables, all model inputs are normalized to ensure consistency. Using the training dataset spanning 1979–2019, we compute the mean and standard deviation for each variable. Normalization is then performed by subtracting the respective mean and dividing by the corresponding standard deviation.
Implementation Details
The main structure of this work follows the backbone of Flash Attention (Dao et al. 2022). We apply the AdamW optimizer with 0.0002 learning rate to the model training. In both global and regional forecasting tasks, we train 100 epochs and set batch size to 16. Our model is trained with PyTorch using 16 NVIDIA Tesla A100 GPUs. In the global forecasting stage, we train the whole model. While in the regional forecasting stage, we only need to train the Spatial-Aligned Attention(SAA) module and freeze the main structure. More implementation details are provided in the logs.
Code Available
We provide some code in the Supplementary Materials for our STCast, OneForecast, GraphCast and related baselines. Baselines are collected from its respective official GitHub repository.
Additional Related Works
Time-series Methods
Before the emergence of recent data-driven forecasting methods on large-scale atmospheric datasets such as ERA5 (Hersbach et al. 2020), several time-series approaches had already been applied to weather forecasting. These earlier works typically framed regional forecasting as a video prediction task, employing spatio-temporal convolutional layers or Transformer blocks to process the input data. For instance, SimVP (Gao et al. 2022) and PastNet (Wu et al. 2024b) employed spatio-temporal convolutions to forecast regional atmospheric images, while STTN (Ji et al. 2024) and PKD-STTN (He, Ji, and Lei 2024) utilized spatio-temporal Transformer blocks. Concurrently, HiSTGNN (Ma et al. 2023) and STCWF (Gong et al. 2024) introduced spatio-temporal graph neural networks and contrastive learning, respectively, to the weather forecasting domain. Notably, PastNet (Wu et al. 2024b) further distinguished itself by incorporating physical principles into its neural network architecture.
Although the aforementioned time-series forecasting methods are commonly referred to as spatio-temporal approaches utilizing spatio-temporal neural networks (NNs), our proposed STCast fundamentally differs from them in several key aspects: (1) Motivation: Previous works aim to capture implicit temporal correlations among time-series inputs and spatial dependencies within the input domain using neural components such as convolutional layers. In contrast, STCast explicitly models temporal correlations across monthly atmospheric patterns and geographical relationships across the global domain. (2) Architecture: Prior works typically model spatial and temporal dimensions through convolution layers or Transformer blocks, where spatial modeling is performed via convolution kernels applied to image grids, and temporal modeling is achieved by extending these kernels across time steps. In contrast, STCast introduces a novel spatio-temporal modeling framework based on monthly Gaussian distributions and global–regional representations, enabling more structured and interpretable learning across both spatial and temporal domains. (3) Benchmarks: Unlike previous time-series methods that are typically evaluated on small-scale, low-resolution datasets such as SEVIR (Veillette, Samsi, and Mattioli 2020) and WD (Wang et al. 2019), STCast is the first to explore explicit spatio-temporal correlations within a realistic Earth system. It is evaluated on the ERA5 dataset (Hersbach et al. 2020), demonstrating its scalability and effectiveness in large-scale global weather forecasting.
| Name | Description | Levels | Resolution | Lat-Lon Range | Time |
| Low-resolution Global Weather Forecasting | |||||
| Z | Geopotential | 13 | -90°S180°W-90°N180°E | 1979-2020 | |
| Q | Specific humidity | 13 | -90°S180°W-90°N180°E | 1979-2020 | |
| U | x-direction wind | 13 | -90°S180°W-90°N180°E | 1979-2020 | |
| V | y-direction wind | 13 | -90°S180°W-90°N180°E | 1979-2020 | |
| T | Temperature | 13 | -90°S180°W-90°N180°E | 1979-2020 | |
| t2m | Temperature at 2m height | Single | -90°S180°W-90°N180°E | 1979-2020 | |
| u10 | x-direction wind at 10m height | Single | -90°S180°W-90°N180°E | 1979-2020 | |
| v10 | y-direction wind at 10m height | Single | -90°S180°W-90°N180°E | 1979-2020 | |
| msl | Mean sea-level pressure | Single | -90°S180°W-90°N180°E | 1979-2020 | |
| sp | Surface pressure | Single | -90°S180°W-90°N180°E | 1979-2020 | |
| High-resolution Regional Weather Forecasting | |||||
| t2m | Temperature at 2m height | Single | 7.5°S114°W-36°N172.5°E | 1979-2020 | |
| u10 | x-direction wind at 10m height | Single | 7.5°S114°W-36°N172.5°E | 1979-2020 | |
| v10 | y-direction wind at 10m height | Single | 7.5°S114°W-36°N172.5°E | 1979-2020 | |
| msl | Mean sea-level pressure | Single | 7.5°S114°W-36°N172.5°E | 1979-2020 | |
| sp | Surface pressure | Single | 7.5°S114°W-36°N172.5°E | 1979-2020 | |
Model Details
The global weather forecasting framework comprises three core components: Encoder, Processor, and Decoder, and two supplementary settings: Reconstruction and Loss Function. In this work, the Encoder and Decoder implementations follow Flash-Attention (Dao et al. 2022), while the Reconstruction setting and Loss Function design adopt the same setting of VA-MoE (Chen et al. 2025). For the forecasting task, the AI model predicts future atmospheric states from historical fields as . Detailed configurations for all five elements are provided below.
Encoder
Atmospheric variables across pressure levels are organized into a 3D tensor , where and denote the global grid height and width, respectively, and is the number of variables. This tensor is projected into patch embeddings via a convolutional layer Conv with stride equal to the patch size:
where we set .
In addition to patch embedding, we incorporate a learnable positional embedding matrix, denoted as , to encode spatial information within the Encoder module. This matrix is initialized using a truncated normal distribution defined as:
where and represent the mean and standard deviation, and and denote the lower and upper bounds of the distribution, respectively. The positional embedding is subsequently added to the input features prior to processing, yielding the updated representation .
Processor
The processor consists of a sequence of Transformer blocks, each comprising multi-head attention, a TMoE (Temporal Mixture-of-Experts) module, layer normalization, and residual connections. The operations within a single block can be formally expressed as:
where Attention denotes the attention mechanism, LN represents layer normalization, and TMoE refers to the Temporal Mixture-of-Experts module as described in the main paper.
Following the design principles of FlashAttention (Dao et al. 2022) and VA-MoE (Chen et al. 2025), we adopt an alternating strategy that combines window-based attention and global self-attention. This hybrid approach enables the model to effectively capture both local and global dependencies in the input distribution.
Decoder
The Decoder module in this work is implemented as a multi-layer perceptron (MLP), which predicts the atmospheric variables for the next timestep. The decoding operation is defined as:
The MLP consists of two linear layers separated by a non-linear activation function. Specifically, the decoding process can be expressed as:
where GELU denotes the Gaussian Error Linear Unit activation function. This structure enables the decoder to model complex relationships in the input features and generate accurate predictions for the subsequent timestep.
Reconstruction
To ensure training stability, we introduce an auxiliary reconstruction path that directly connects the Encoder and Decoder modules to reconstruct the input variables. This design complements the primary prediction path, which consists of the Encoder, Processor, and Decoder modules. Notably, the Encoder and Decoder are shared across both paths. The overall process is defined as:
where Encoder, Processor, and Decoder denote the respective modules in our framework. The reconstruction path facilitates the learning of robust representations by encouraging the model to preserve essential input information throughout the encoding and decoding stages.
Under this configuration, the Encoder and Decoder modules are dedicated to encoding and decoding the input variables, respectively, while the Processor is solely responsible for prediction. By excluding the Processor from the encoding and decoding stages, the framework avoids unnecessary computational overhead, thereby enhancing efficiency without compromising performance.
Loss function
To address both prediction and reconstruction tasks, we employ the L2 loss function to quantify point-wise errors between the predicted outputs and the ground truth. The prediction loss and reconstruction loss are defined as follows:
where and denote the reconstructed and predicted outputs, respectively. The operator Mean computes the average error across multiple dimensions. A weighting hyperparameter is introduced to balance the two objectives in the final loss function.
| Model | Params(M) | MACs(G) | GPUs | Training Time |
| Fengwu (Chen et al. 2023a) | 153.49 | 132.83 | 32 A100 | 17 days |
| FourCastNet (Kurth et al. 2023) | - | - | 64 A100 | 16 hrs |
| GraphCast (Lam et al. 2023) | 28.95 | 1639.26 | 32 TPUv4 | 4 weeks |
| Pangu-Weather (Bi et al. 2023) | 23.83 | 142.39 | 192 V100 | 64 days |
| VA-MoE (Chen et al. 2025) | 665.37 | - | 32 A100 | 6 days |
| OneForecast (Gao et al. 2025) | 24.76 | 509.27 | 16 A100∗ | 8 days∗ |
| Ours (STCast) | 616.16 | 436.12 | 16 A100 | 5 days |
Experiments Details
Evaluation Metric
In this work, we evaluate the forecasting performance between our STCast and other methods on RMSE (Root Mean Square Error) and ACC (Anomalous Correlation Coefficient), which can be defined as:
where and denotes the predicted variables and ground-truth at the horizontal coordinate and time ; and denote the length of latitude and longitude in the global region.
Considering the difference in the distribution of atmospheric variables at latitudes, we introduce the latitude-dependent function to weight the atmospheric variables. The function is formulated as:
where and denote the latitude at index and , respectively.
For typhoon track prediction, we evaluate model performance using two metrics: Mean Distance Error (MDE) and Haversine Distance. First, the Haversine Distance is computed between the predicted typhoon center and the ground-truth location to account for the curvature of the Earth. Subsequently, the MDE is used to quantify the average positional error across all time steps. These metrics are defined as follows:
| MDE | (19) | |||
where is the average radius of the earth, and are the latitude and longitude of two points in earth system. and are the latitude-longitude coordinates of predicted and real points.
Typhoon Tracking
Following previous AI-based approaches (Bi et al. 2023; Magnusson et al. 2021), we identify the eye of a tropical cyclone as the location of the local minimum in mean sea level pressure (MSLP). For tracking purposes, the temporal resolution is set to 6-hour intervals.
This study focuses on two extreme cyclones: Typhoon Ewiniar and Typhoon Yinxing. Typhoon Ewiniar formed east of Mindanao on May 24, 2024, traversed the Philippine Sea, and recurved northeastward over the Okinawa–Ryukyu region. Typhoon Yinxing developed east of Yap Island on November 4, 2024, crossed the Philippine Sea, and entered the South China Sea. The initial conditions for these cyclones are set at 00:00 UTC on May 24, 2024, and 00:00 UTC on November 4, 2024, respectively. Ground-truth and ECMWF are obtained from TCData222tcdata.typhoon.org.cn.
Additional Results
Efficiency Analysis
As shown in Tab. 5, we compare the number of parameters, multiply–accumulate operations (MACs), GPU usage, and training duration across six baseline models. Although STCast contains more parameters and MACs than GNN-based methods such as GraphCast and OneForecast, its overall computational cost remains significantly lower than that of previous models, particularly Fengwu, Pangu-Weather, and GraphCast. These results demonstrate that STCast achieves superior performance while maintaining comparable computational efficiency.
Additional Global Forecasting Analysis
As illustrated in Fig. 10, we compare STCast with several state-of-the-art models across forecasting horizons ranging from 6 to 120 hours, evaluated on four key atmospheric variables. Experimental results show that STCast performs comparably to VA-MoE and Stormer in predicting Z500 and U10, while outperforming all baselines in T850 and T2M. These findings highlight the effectiveness of STCast and its integrated TMoE architecture in global weather forecasting, demonstrating its ability to capture both temporal correlations and seasonal variability across diverse meteorological variables.
Additional Typhoon Track Prediction
In addition to the typhoon track visualizations presented in the main paper, we compare the Mean Distance Error (MDE, in kilometers) between ECMWF and STCast, as shown in Fig. 11. Experimental results indicate that STCast achieves comparable performance to ECMWF in short-term forecasts, while significantly outperforming it in long-term predictions. Specifically, STCast yields track prediction errors of 67.10 km for Typhoon Yinxing and 109.34 km for Typhoon Ewiniar, substantially lower than ECMWF’s errors of 165.25 km and 138.82 km, respectively. These findings demonstrate the strong capability of STCast in extreme event assessment, particularly in accurately forecasting tropical cyclone trajectories over extended time horizons.
Additional Visualization
We provide more visualization about regional weather forecasting of 6-hour, 0.5-day, 1-day, 1.5-day, 2-day, 2.5-day, 3-day, 3.5-day, 4-day, 4.5-day, 5-day, 5.5-day, 6-day, 6.5-day, 7-day, 7.5-day, 8-day, 8.5-day, 9-day, 9.5-day, and 10-day in Fig. 12, Fig. 13, Fig. 14, Fig. 15, Fig. 16, Fig. 17, Fig. 18, Fig. 19, Fig. 20, Fig. 21, Fig. 22, Fig. 23, Fig. 24, Fig. 25, Fig. 26, Fig. 27, Fig. 28, Fig. 29, Fig. 30, Fig. 31, and Fig. 32.
We also provide more visualization about global weather forecasting of 6-hour, 0.5-day, 1-day, 1.5-day, 2-day, 2.5-day, 3-day, 3.5-day, 4-day, 4.5-day, 5-day, 5.5-day, 6-day, 6.5-day, 7-day, 7.5-day, 8-day, 8.5-day, 9-day, 9.5-day, and 10-day in Fig. 33, Fig. 34, Fig. 35, Fig. 36, Fig. 37, Fig. 38, Fig. 39, Fig. 40, Fig. 41, Fig. 42, Fig. 43, Fig. 44, Fig. 45, Fig. 46, Fig. 47, Fig. 48, Fig. 49, Fig. 50, Fig. 51, Fig. 52, and Fig. 53.